OpenStudy (anonymous):

-4c-c+12=-16+3 plz tell

OpenStudy (anonymous):

plz show work

myininaya (myininaya):

thats an equation!

OpenStudy (anonymous):

ok now whats the anser

myininaya (myininaya):

i don't feel like answering your question since you like to argue bye! goodluck!

OpenStudy (anonymous):

-4c-c+12=-13 -4c-c==-25 -5c=-25 c=5

OpenStudy (anonymous):

si correcto but next time ask politely when u need an answer